# CausalDynamics: A large-scale benchmark for structural discovery of dynamical causal models

### v1

@@ Abstract
Causal discovery for dynamical systems poses a major challenge in fields where active interventions are infeasible. Most methods used to investigate these systems and their associated benchmarks are tailored to deterministic, low-dimensional and weakly nonlinear time-series data. To address these limitations, we present CausalDynamics, a large-scale benchmark and extensible data generation framework to advance the structural discovery of dynamical causal models. Our benchmark consists of true causal graphs derived from thousands of coupled ordinary and stochastic differential equations as well as two idealized climate models. We perform a comprehensive evaluation of state-of-the-art causal discovery algorithms for graph reconstruction on systems with noisy, confounded, and lagged dynamics. CausalDynamics consists of a plug-and-play, build-your-own coupling workflow that enables the construction of a hierarchy of physical systems. We anticipate that our framework will facilitate the development of robust causal discovery algorithms that are broadly applicable across domains while addressing their unique challenges. We provide a user-friendly implementation and documentation on https://kausable.github.io/CausalDynamics.
@@
1 Introduction
Understanding causal mechanisms in nonlinear dynamical systems is crucial across many fields. However, direct interventions are often impractical or impossible. An alternate strategy for encoding causal information is through detailed physical modeling, such as Earth System models, which relies on ad-hoc parameterizations and can be computationally expensive. In contrast, data-driven causal discovery frameworks have emerged as a promising avenue to infer cause–effect relationships directly from time series observations, obviating the need for interventions or simulations. Originating from Wright’s path analysis in the 1930s [1] and later formalized by Wiener [2] and Granger [3] in ${}^{\text{th}}$ century, causal inference now includes deep learning (DL)-based methods [4, 5], designed to infer complex, nonlinear, and stochastic causal structures.
@@ 1 Introduction
Despite rapid methodological advances, there exists no standardized benchmark for evaluating causal discovery in highly nonlinear, dynamical settings [6, 7, 8], which is critical to understand and predict the behavior of physical systems [9, 10, 11, 12]. Most benchmarks are built on synthetic data generated by static causal graphs (e.g., [13]) or auto-regressive models (e.g., [14, 15]) for domain-specific applications [16, 17]. Limited examples from the real-world are available but lack a fully resolved causal ground truth [18, 19, 20]. As a result, most methods are validated on toy systems or on real-world data failing to capture continuous state-space developments, complex feedback loops, stochasticity, and regime shifts, making it impossible to isolate algorithmic limitations (e.g., varsortability [21, 22]) from dataset characteristics.
We posit that advancing the development of new causal discovery methods capable of capturing high dimensional, nonlinear, time-lagged, and noisy structures with unobserved confounding, often found in real-world physical systems [23], requires a novel benchmark. In fact, examples from various fields demonstrate how such benchmarks also play a vital role in unlocking methodological innovation for data-driven methods and boosting scientific progress. A prominent example is CASP13, a protein structure modeling benchmark that enabled the development of AlphaFold, which was awarded the 2024 Nobel Prize in Chemistry [24, 25, 26]. Other benchmarks include ImageNet [27] which highlighted the computer vision capabilities of AlexNet [28] or NWPU-RESISC45 which first demonstrated the superior skill of deep learning approaches for remote-sensing classification [29].
In this paper, we present CausalDynamics, a tiered benchmark and extensible data generation framework, as shown in Figure 1, to advance the structural discovery of dynamical causal models. The simple tier contains true causal graphs for hundreds of three-dimensional chaotic dynamical systems. In the second tier, we adapt an existing algorithm [30] for graph generation to couple deterministic and stochastic dynamical systems with periodic functions to construct thousands of complex graph structures. This framework allows us to address the set of challenges often found in real-world systems as highlighted above. The final tier contains the true causal graphs for two idealized atmosphere-ocean models including multiple coupling experiments for different modes of variability.
@@
  • CausalDynamics, the largest benchmark for causal discovery methods for stochastic chaotic systems containing over 5300 graphs and over 50 million preprocessed samples.
@@ 3.1 Benchmarks for causal discovery
Real-world. Real-world datasets for causal discovery remain rare because true causal graphs are difficult to obtain. A handful of benchmarks provides time series data, e.g. human-motion capture (MoCap) [5], S&P 100 stock prices [20], climate variability and teleconnections [43], river discharge (CausalRivers) [18], and ecological observations [44]. However, most observed collections focus on static, domain-specific graphs, such as large-scale single-cell RNA perturbations (CausalBench) [16, 17], the Old Faithful geyser eruptions [45] and immune-cell protein networks [46, 47]. Bivariate settings like the Tübingen Cause Effect Pairs [34] further complement these resources.
Synthetic. Due to the domain specifications, limited availability and often low dimensionality of real-world datasets, synthetic data plays an important role in understanding complex systems in general [48, 49, 50, 51] and benchmarking causal discovery methods [6] specifically. In static settings, SCMs or structural vector autoregressive (SVAR) models sample causal coefficients and noise to generate graphs [22] often following multivariate linear or nonlinear functions [45, 52, 53], or physical laws [54]. More realistic pipelines derive the causal graph to fit coefficients from observational data (e.g. causalAssembly [13], SynTReN [55]), which limits their application to the domain and the properties of the underlying observed dataset. For time series, idealized dynamical systems, e.g., Lorenz or Rössler Oscillators [56], underpin simpler benchmarks [5, 57, 58, 59], while domain-informed models include fMRI networks (Netsim) [60], gene regulatory networks (DREAM3 & 4) [61, 62] or financial data (FINANCE) [63, 64]. A collection of climate and weather benchmarks can be found on CauseMe (https://causeme.uv.es/) [19, 65] with pseudo-realistic climate data generated using the SAVAR model [15]. Going beyond single-domain applications, novel approaches propose pipelines to flexibly generate time series datasets for a range of properties and natural systems [66] from any observational dataset using DL (e.g., CausalTime [14]).
@@ 3.2 Causal discovery methods
(i) Granger causality (GC) [3] is one of the oldest concepts in causal inference [67]. GC tests whether a given effect is optimally forecast by its causes under the assumptions of no unobserved confounders. GC often fails for nonlinear dynamics [44], leading to deep-learning variants: Neural GC (NGC) [5] employs regularized multilayer perceptrons (cMLP) or long short-term memory (cLSTM) to learn nonlinear autoregressive links, while CUTS+ uses a coarse-to-fine pipeline with a message-passing graph neural network to recover causal graphs from high-dimensional, irregular time series [68].
(ii) Constraint-based methods infer causal structure by enforcing the statistical constraints implied by conditional independencies in the data (e.g., PC [37]). PCMCI+ (Peter Clark Momentary Conditional Independence) [40] adapts PC to multivariate time series by preselecting candidate parents and performing Momentary Conditional Independence (MCI) tests over all time-lags to infer contemporaneous and lagged edges. F-PCMCI boosts scalability by prefiltering parents via transfer-entropy before MCI testing [69].
(iii) Noise-based causal discovery methods exploit the fact that in a correctly specified SCM, the noise term is statistically independent of its inputs only in the true causal direction. For example, LiNGAM [70] and its temporal extension VARLiNGAM [71] assume a linear model with non-Gaussian noise and recover a unique DAG and lagged links by analysis the residuals, while additive noise models (ANMs) (e.g., [45]) generalize this to nonlinear relations by choosing the direction where residuals remain independent of their input.
(iv) Score-based learning algorithms infer causal relationships between variables by evaluating and ranking potential causal graphs based on a score function such as least-squares error. For instance, DYNOTEARS [20] adopts a score-based SVAR formulation with a penalized least-squares loss and a differentiable acyclicity constraint to jointly recover instantaneous and lagged weights.
(v) Topology-based approaches are based on Takens’ theorem [72] to reconstruct the attractor of a dynamical system from delay-embedded time series data. TSCI (Tangent Space Causal Inference) [73] then applies Convergent Cross Mapping [44] within each manifold’s tangent space to test for causal influence by assessing how well the state-space of one variable predicts another.
@@ 4.1 Tier 1 – Simple causal models
In the simple complexity tier, we select 59 three-dimensional systems from the dysts benchmark [56, 74] and derive their SDCMs as adjacency matrices. For each system, we simulate trajectories with 10 random initial conditions over 1000 time steps.
@@ 4.2 Tier 2 – Hierarchically coupled causal models
$n\in\mathbb{N},r\in[0,1]$
2:$\mathcal{A}\leftarrow\text{zeros}(n\times n)$; $K\leftarrow\text{zeros}(n)$; $v_{a}\leftarrow\text{zeros}(n)$triangleright adj. matrix; attachm. kernel; ancestors
@@
7:$K[0]\leftarrow 1$; $v_{a}[0]\leftarrow 0$triangleright Initialize attachment kernel and ancestors
8:for $i\leftarrow 1$ to $n-1$ do
9:     $K^{\text{prob}}\leftarrow K[0:i]/\sum(K[0:i])$triangleright Normalize to get probabilities
10:     $m\leftarrow\text{Multinomial}(K^{\text{prob}},1)$triangleright Sample node based on probabilities
@@
12:         $m\leftarrow v_{a}[m]$triangleright Redirect to ancestor node
@@
14:     $\mathcal{A}[i,m]\leftarrow 1$triangleright Add edge in graph
15:     $K[i]\leftarrow K[i]+1$; $K[m]\leftarrow K[m]+1$triangleright Update source & target attachment kernels
16:     $v_{a}[i]\leftarrow m$triangleright Track ancestor
@@
Edges. Non-root nodes integrate the transformed signals received from their parent nodes via the information carried by the edges [76, 77]. Each edge $(k,i)\in\mathcal{E}$ is realized as an MLP [78] without activation:
$f_{(k,i)}(x_{k}(t))=W_{(k,i)}x_{k}(t)+b_{(k,i)},\quad W_{(k,i)}\sim\mathcal{N}(0,1),\quad b_{(k,i)}\sim\mathcal{N}(0,1),$ (4)
where $W_{(k,i)}\in\mathbb{R}^{d\times d}$. Weights are sparsified with a dropout probability $p_{\text{zero}}$. The value of node $v_{k}$ at time $t$ is computed by aggregating its incoming signals:
@@
Figure 3: (a) Hierarchically coupled graphs are sample as scale-free DAGs, where root nodes can either be driven by period functions (root node 1) or dynamical systems (root node 2). Information is passed to the leaf nodes (leaf node 0) through MLPs. To correct for varsortability node values can optionally be standardized. To create a diverse set of challenges we introduce (b) unobserved confounders (green dotted lines), or (c) time-lag $\tau$ (red node). Assigned drivers are plotted by solid black edges. Note that each node is $\in\mathbb{R}^{d}$ and each edge is a function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$.
@@ 4.3 Tier 3 – Pseudo-real physical systems
In a low-dimensional setting, we model the El Niño–Southern Oscillation (ENSO) as implemented in the XRO package [83], which merges the Hasselmann stochastic framework [84] and recharge oscillator dynamics [85]. As XRO initializes with observed sea surface temperatures and thermocline depth, it reproduces key observational ENSO statistics [86]. The package also links multiple ocean-basin modes, allowing for adjustable coupling strengths to simulate tightly coupled or largely independent behavior. To simulate higher dimensional atmosphere-ocean dynamics, we utilize the qgs package of a simplified quasi-geostrophic two-layer model resolving barotropic and baroclinic interactions derived from the Modular Arbitrary-Order Ocean-Atmosphere Model (MAOOAM) [87].
@@ 4.4 Dataset summary
We provide a summary of the preprocessed dataset in Table 1 and Figure 13. Unless otherwise stated, each graph constitutes 10 randomly initialized trajectories with 1000 time steps. In total, we generated 585 simple, 4745 coupled, and 12 climate graphs, for a total of 5342 graphs. More details on the selected parameters can be found in Appendix D.
@@
Coupled Coupled ODE/SDE ($N{=}\{3,5,10\}$) Confounder, time-lag, standardized, forcing 4745
@@ 4.5 Evaluation metrics
Our benchmark is concerned with reconstructing the true causal graph of the underlying dynamical systems. To evaluate the similarity of the graphs inferred by the baselines and the true causal graph we estimate Area Under the Receiver Operating Characteristic (AUROC) [88, 89] and Area Under the Precision Recall Curve (AUPRC) [90] scores by comparing the true and predicted adjacency matrices. Our full evaluation workflow including details on baseline methods is provided in Appendix E.

### v2

# CausalDynamics: A large‐scale benchmark for structural discovery of dynamical causal models

@@ Abstract
Causal discovery for dynamical systems poses a major challenge in fields where active interventions are infeasible. Most methods used to investigate these systems and their associated benchmarks are tailored to deterministic, low-dimensional and weakly nonlinear time-series data. To address these limitations, we present CausalDynamics, a large-scale benchmark and extensible data generation framework to advance the structural discovery of dynamical causal models. Our benchmark consists of true causal graphs derived from thousands of both linearly and nonlinearly coupled ordinary and stochastic differential equations as well as two idealized climate models. We perform a comprehensive evaluation of state-of-the-art causal discovery algorithms for graph reconstruction on systems with noisy, confounded, and lagged dynamics. CausalDynamics consists of a plug-and-play, build-your-own coupling workflow that enables the construction of a hierarchy of physical systems. We anticipate that our framework will facilitate the development of robust causal discovery algorithms that are broadly applicable across domains while addressing their unique challenges. We provide a user-friendly implementation and documentation on https://kausable.github.io/CausalDynamics.
1 Introduction
@@
Despite rapid methodological advances, there exists no standardized benchmark for evaluating causal discovery in highly nonlinear, dynamical settings [6, 7, 8], which is critical to understand and predict the behavior of physical systems [9, 10, 11, 12, 13, 14]. Most benchmarks are built on synthetic data generated by static causal graphs (e.g., [15]) or auto-regressive models (e.g., [16, 17]) for domain-specific applications [18, 19]. Limited examples from the real-world are available but lack a fully resolved causal ground truth [20, 21, 22]. As a result, most methods are validated on toy systems or on real-world data failing to capture continuous state-space developments, complex feedback loops, stochasticity, and regime shifts, making it impossible to isolate algorithmic limitations (e.g., varsortability [23, 24]) from dataset characteristics.
We posit that advancing the development of new causal discovery methods capable of capturing high dimensional, nonlinear, time-lagged, sparse, and noisy structures with unobserved confounding, often found in real-world physical systems [25, 26, 27], requires a novel benchmark. In fact, examples from various fields demonstrate how such benchmarks also play a vital role in unlocking methodological innovation for data-driven methods and boosting scientific progress. A prominent example includes CASP13, a protein structure modeling benchmark that enabled the development of AlphaFold, a co-awardee of the 2024 Nobel Prize in Chemistry [28, 29, 30]. Other benchmarks include ImageNet [31] which highlighted the computer vision capabilities of AlexNet [32] or NWPU-RESISC45 which first demonstrated the superior skill of deep learning approaches for remote-sensing classification [33].
In this paper, we present CausalDynamics, a tiered benchmark and extensible data generation framework, as shown in Figure 1, to advance the structural discovery of dynamical causal models. The simple tier contains true causal graphs for hundreds of three-dimensional chaotic dynamical systems. In the second tier, we adapt an existing algorithm [34] for graph generation to couple deterministic and stochastic dynamical systems with periodic functions for constructing thousands of complex graph structures. This framework allows us to address the set of challenges often found in real-world systems as highlighted previously. The final tier contains the true causal graphs for two idealized atmosphere-ocean models including multiple coupling experiments of different modes of variability.
@@
  • CausalDynamics, the largest benchmark for causal discovery methods for stochastic chaotic systems containing over 14000 graphs and over 50 million preprocessed samples.
@@ 3.1 Benchmarks for causal discovery
Real‑world. Real‑world datasets for causal discovery remain rare because true causal graphs are difficult to obtain. A handful of benchmarks provides time series data, e.g. human-motion capture (MoCap) [5], S&P 100 stock prices [22], climate variability and teleconnections [47], river discharge (CausalRivers) [20], and ecological observations [48]. However, most observed collections focus on static, domain-specific graphs, such as large‑scale single‑cell RNA perturbations (CausalBench) [18, 19], the Old Faithful geyser eruptions [49] and immune‑cell protein networks [50, 51]. Bivariate settings like the Tübingen Cause Effect Pairs [38] further complement these resources.
Synthetic. Due to the domain specifications, limited availability and often low dimensionality of real-world datasets, synthetic data plays an important role in understanding complex systems in general [52, 53, 54, 55] and benchmarking causal discovery methods [6] specifically. In static settings, SCMs or structural vector autoregressive (SVAR) models sample causal coefficients and noise to generate graphs [24] often following multivariate linear or nonlinear functions [49, 56, 57], or physical laws [58]. More realistic pipelines derive the causal graph to fit coefficients from observational data (e.g. causalAssembly [15], SynTReN [59]), which limits their application to the domain and the properties of the underlying observed dataset. For time series, idealized dynamical systems, e.g., Lorenz or Rössler Oscillators [60], underpin simpler benchmarks [5, 61, 62, 63], while domain‑informed models include fMRI networks (Netsim) [64], gene regulatory networks (DREAM3 & 4) [65, 66] or financial data (FINANCE) [67, 68]. A collection of climate and weather benchmarks can be found on CauseMe (https://causeme.uv.es/) [21, 69] with pseudo-realistic climate data generated using the SAVAR model [17]. Going beyond single-domain applications, novel approaches propose pipelines to flexibly generate time series datasets for a range of properties and natural systems [70] from any observational dataset using DL (e.g., CausalTime [16]).
@@ 3.2 Causal discovery methods
(i) Granger causality (GC) [3] is one of the oldest concepts in causal inference [71]. GC tests whether a given effect is optimally forecast by its causes under the assumptions of no unobserved confounders. GC often fails for nonlinear dynamics [48], leading to deep‑learning variants: Neural GC (NGC) [5] employs regularized multilayer perceptrons (cMLP) or long short-term memory (cLSTM) to learn nonlinear autoregressive links, while CUTS+ uses a coarse‑to‑fine pipeline with a message‑passing graph neural network to recover causal graphs from high‑dimensional, irregular time series [72]. Another notable time-series extension is Temporal Causal Discovery Framework (TCDF) [68], which leverages convolutional neural networks with attention mechanisms to identify lagged causal links directly from multivariate sequences.
(ii) Constraint-based methods infer causal structure by enforcing the statistical constraints implied by conditional independencies in the data (e.g., PC [41]). PCMCI+ (Peter Clark Momentary Conditional Independence) [44] adapts PC to multivariate time series by preselecting candidate parents and performing Momentary Conditional Independence (MCI) tests over all time-lags to infer contemporaneous and lagged edges. F‑PCMCI boosts scalability by prefiltering parents via transfer‑entropy before MCI testing [73]. More recently, permutation-based methods such as Greedy Relaxations of the Sparsest Permutation (GRaSP) [74] have been developed, which search over permutations of variables and iteratively prune spurious edges to identify sparse causal graphs.
(iii) Noise-based causal discovery methods exploit the fact that in a correctly specified SCM, the noise term is statistically independent of its inputs only in the true causal direction. For example, LiNGAM [75] and its temporal extension VARLiNGAM [76] assume a linear model with non‑Gaussian noise and recover a unique DAG and lagged links by analysis the residuals, while additive noise models (ANMs) (e.g., [49]) generalize this to nonlinear relations by choosing the direction where residuals remain independent of their input. The recent Repetitive Causal Discovery (RCD) algorithm [77] extends the LiNGAM family by incorporating constrained functional causal models with regularization, enabling more robust identification in high-dimensional or noisy settings.
(iv) Score-based learning algorithms infer causal relationships between variables by evaluating and ranking potential causal graphs based on a score function such as least-squares error. For instance, DYNOTEARS [22] adopts a score-based SVAR formulation with a penalized least‑squares loss and a differentiable acyclicity constraint to jointly recover instantaneous and lagged weights.
(v) Topology-based approaches are based on Takens’ theorem [78] to reconstruct the attractor of a dynamical system from delay‐embedded time series data. TSCI (Tangent Space Causal Inference) [79] then applies Convergent Cross Mapping [48] within each manifold’s tangent space to test for causal influence by assessing how well the state-space of one variable predicts another.
@@ 4.1 Tier 1 – Simple causal models
In the simple complexity tier, we select 59 three‐dimensional systems from the dysts benchmark [60, 80] and derive their SDCMs as adjacency matrices. For each system, we simulate trajectories with 5 random initial conditions over 1000 time steps.
@@ 4.2 Tier 2 – Hierarchically coupled causal models
$1]n\in\mathbb{N},r\in[0,$1]
[…]
@@
[…]
8:for $i\leftarrow 1$ to n1n-1 do
9:  $K^{\text{prob}}\leftarrow K[0:i]/\sum(K[0:i])$ $\triangleright$ Normalize to get probabilities
10:  $m\leftarrow\text{Multinomial}(K^{\text{prob}},1)$ $\triangleright$ Sample node based on probabilities
@@
12:   $m\leftarrow v_{a}[m]$ $\triangleright$ Redirect to ancestor node
@@
14:  $\mathcal{A}[i,m]\leftarrow 1$ $\triangleright$ Add edge in graph
15:  $K[i]\leftarrow K[i]+1$; $K[m]\leftarrow K[m]+1$ $\triangleright$ Update source & target attachment kernels
16:  $v_{a}[i]\leftarrow m$ $\triangleright$ Track ancestor
@@
  • Linear drivers: Linear functions $x_{v_{k}}(t)=mt+b$} modeling continuous temporal changes such as rising global mean temperatures in first order approximation.
Edges. Non‑root nodes integrate the transformed signals received from their parent nodes via the information carried by the edges [82, 83]. Each edge $(k,i)\in\mathcal{E}$ is realized as an MLP [84] with optional activations $\phi_{(k,i)}$:
$f_{(k,i)}(x_{k}(t))$ $=\phi_{(k,i)}\left(W_{(k,i)}x_{k}(t)+b_{(k,i)}\right),$ (4)
$W_{(k,i)}$ $\sim\mathcal{N}(0,1),~b_{(k,i)}\sim\mathcal{N}(0,1),$
$\phi_{(k,i)}$ $\sim\text{Uniform}\left(\mathcal{\{\text{identity},\text{sin},\text{sigmoid},\text{tanh},\text{ReLU}\}}\right)$
where $W_{(k,i)}\in\mathbb{R}^{d\times d}$ and $\phi_{(k,i)}=\text{identity}$ in case of no activations. Weights are sparsified with a dropout probability $p_{\text{zero}}$. The value of node $v_{k}$ at time $t$ is computed by aggregating its incoming signals:
@@
Figure 3: (a) Hierarchically coupled graphs are sample as scale-free DAGs, where root nodes can either be driven by period functions (root node 1) or dynamical systems (root node 2). Information is passed to the leaf nodes (leaf node 0) through MLPs. To correct for varsortability node values can optionally be standardized. To create a diverse set of challenges we introduce (b) unobserved confounders (green dotted lines), or (c) time-lag $\tau$ (red node). The type of nonlinearity can also be varied by prescribing or randomly assigning different edge-level activation functions. Assigned drivers are plotted by solid black edges. Note that each node is $\in\mathbb{R}^{d}$ and each edge is a function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$.
@@ 4.3 Tier 3 – Pseudo-real physical systems
In a low-dimensional setting, we model the El Niño–Southern Oscillation (ENSO) as implemented in the XRO package [89], which merges the Hasselmann stochastic framework [90] and recharge oscillator dynamics [91]. As XRO initializes with observed sea surface temperatures and thermocline depth, it reproduces key observational ENSO statistics [92]. The package also links multiple ocean‐basin modes, allowing for adjustable coupling strengths to simulate tightly coupled or largely independent behavior. To simulate higher dimensional atmosphere-ocean dynamics, we utilize the qgs package of a simplified quasi-geostrophic two-layer model resolving barotropic and baroclinic interactions derived from the Modular Arbitrary-Order Ocean-Atmosphere Model (MAOOAM) [93].
@@ 4.4 Dataset summary
We provide a summary of the preprocessed dataset in Table 1 and Figure 13. Unless otherwise stated, each graph constitutes 5 randomly initialized trajectories with 1000 time steps. In total, we generated 585 simple, 14096 coupled, and 12 climate graphs, for a total of 14693 graphs. More details on the selected parameters can be found in Appendix D.
@@
Coupled Coupled ODE/SDE ($N{=}\{3,5,10\}$) Confounder, time-lag, standardized, forcing 14096
@@ 4.5 Evaluation metrics
Our benchmark is concerned with reconstructing the true causal graph of the underlying dynamical systems. To evaluate the similarity of the graphs inferred by the baselines and the true causal graph we estimate Area Under the Receiver Operating Characteristic (AUROC) [94, 95], Area Under the Precision Recall Curve (AUPRC) [96], and Structural Hamming Distance (SHD) [97, 98] scores by comparing the true and predicted adjacency matrices. Our full evaluation workflow including details on baseline methods is provided in Appendix E.